### v4

# The Less Intelligent the Elements, the More Intelligent the Whole. Or, Possibly Not?

Abstract
We approach the debate on how “intelligent” artificial agents should be, by endowing the preys and predators of the Lotka-Volterra model with behavioural algorithms characterized by different levels of sophistication. We find that by endowing both preys and predators with the capability of making predictions based on linear extrapolation a novel sort of dynamic equilibrium appears, where both species co-exist while both populations grow indefinitely. While we confirm that, in general, simple agents favour the emergence of complex collective behaviour, we also suggest that the capability of individuals to take first-order derivatives of one other’s behaviour may allow the collective computation of derivatives of any order.
Keywords: Swarm Intelligence, Collective Intelligence, Crowd Wisdom, Social Connectionism, Prey-Predator, Unlimited Growth
1 Introduction
Agent-Based Models (ABMs) have been introduced into the social sciences with the aim of reproducing the emergence of macroscopic regularities out of microscopic interactions between relatively simple artificial agents [26], a promise that has obtained some success amid slow but steadily growing diffusion of ABMs in economics, sociology, political science and etology [54]. This paradigm has conceived ABMs as sufficiency proofs that would identify the minimalist microscopic conditions for specific macroscopic dynamics to emerge [25], an attitude succintly described by the KISS principle (Keep It Simple, Stupid) [22]. This posture resonates with the connectionist philosophy of obtaining complex behaviour out of interactions between relatively simple elements, to be applied this time to interacting humans instead of interacting neurons [93] [18].
However, growing diffusion and realism of ABMs eventually suggested that artificial agents should not be too simple in order to obtain meaningful results. The alternative KIDS principle (Keep It Detailed, Stupid) suggests that artificial agents should be endowed with fairly sophisticated capabilities in order to generate relevant macroscopic regularities [21]. Thus the question is, how “intelligent” should artificial agents be? Do criteria exist, that allow modelers to estimate how detailed their artificial agents should be, in order to generate meaningful emergent properties? Some attempts have been made at resolving this debate by comparing specific ABMs, but results have been largely inconclusive hitherto [20] [79].
We approached this question from a slightly different point of view. We focused on one single model, namely an ABM version of the Lotka-Volterra prey-predator model [52] [90]. The reason to select this model is that, beyond obvious ecological interpretations, it is applied to social dynamics as well [16] [94] [101] [17] [59] [40] [92]. Within its framework we explored forms of individual intelligence that have been suggested by empirical findings, theoretical statements and numerical simulations in a wide array of disciplines to be possibly conducive to collective intelligence.
The logic of our method mirrors the practice of exploring the collective consequences of individual behavior in the simulated and iterated multi-player Prisoner’s Dilemma [6] [56]. Just like researchers ask what individual behaviors generate cooperation in the Prisoner’s Dilemma, we ask what individual behaviors generate co-existence of predators and preys in the Lotka-Volterra model.
The rest of this article is organized as follows. In the ensuing section § 2 we distill propositions from several disciplines concerning the relations between individual and collective intelligence. In § 3 we translate these propositions into behavioral algorithms that we ascribe to either predators or preys. Unexpectedly, besides behaviors that generate co-existence of stable populations we also found behaviors that generate co-existence of exploding populations of both predators and preys, a circumstance that we discuss in the concluding section § 4 as expressing unlimited growth in capitalistic economies. The mathematical properties of the Lotka-Volterra model, the parameters and outputs of our model and a sensitivity analysis of its results are expounded in appendices § A, § B and § C, respectively. The code of our extended Lotka-Volterra model is available on CoMSES.
2 Individual and Collective Intelligence
In this section we review theoretical insights and empirical evidence that bear on the question whether a collectivity works better if their members understand its underlying mechanisms, or rather base their decisions on simple rules. We subsumed our findings into three propositions, the third of which comes in three versions. We list these propositions in ascending order, from KISS- to increasingly KIDS-supportive.
Furthermore, we also collected three technical qualifications that may be relevant to these propositions. These technical qualifications are expounded in § 2.1.
Our first proposition captures positions that justify the KISS principle on theoretical and empirical reasons. According to their proponents, KISS is not just a convenient and sometimes sufficient assumption but also what humans actually do (descriptive value), as well as what they should do (normative value).
One extreme version maintains that individual intelligence is totally irrelevant for social science [65]. Inparticular, several “zero-intelligence” models can be found in economics, notably for double-auction markets [33] [34], financial markets [23] and, more in general, any sort of prediction-based markets [67]. Such models typically suggest that markets — even futures markets — can work perfectly well with decision makers who have no memory and are incapable of cognition of whatsoever form. However, some of these models have been found to hide constraints that artificially generate their most impressive results [32].
A less extreme version identifies Fast and Frugal Heuristics (FFH) that are reasonably successful, robust, and computationally cheap [31] [69][64] [28] [27]. This research stream stresses that simple heuristics are in general more effective than complex optimization procedures.
The FFH paradigm does not understand bounded rationality [76] [77] as a deviation from optimality but rather as the evolutionary optimal answer to radical uncertainty and imperfect information [71] [39] [30] [63]. Consequently, FFH have descriptive as well as prescriptive value [29] [43] [37] [19].
Similar conclusions are suggested by The Traveler’s Dilemma, a game where rational players endowed with unlimited ability to read one other’s minds (“I think that she thinks that I think that …”) receive lower payoffs than boundedly rational players who carry out mind reading for a couple of steps at most [8] [9]. Interestingly, when this game is played with real people they invariably behave the “stupid,” more efficient way.
Finally, ABMs of the stock market where both sophisticated and simple agents operate have generated the unexpected finding that sophisticated agents typically make lower profits than the simple ones [82] [36]. In this specific case, the reason was that sophisticated estimation techniques suffered from overfitting a noisy environment. By contrast, simple rules traded-off their imperfections with their robustness.
Let us condense these KISS-supportive experiences into the following proposition:
Proposition 1.
There exist settings where lack of individual intelligence is necessary to reach collective intelligence.
Let us now move onto a set of insights that focus on a very special aspect of individual intelligence, namely the ability to make predictions. Prediction can be very local and very limited, such as linear extrapolation, or it may involve extremely sophisticated cognitive abilities. Depending on flavour, it can span the whole gamut from moderate KISS to maximum KIDS. In this section we limit ourselves to enunciate a general principle that we shall decline into specific choices in § 3.
Several ABMs have highlighted the importance of predictions. For instance, it has been found that robots that are capable of mapping their environment and predicting what other robots will eventually do perform better than those that are based on FFH [73]. Likewise, models of distributed resource allocation have highlighted that agents that are able to extrapolate general trends perform better than those who do not [45] [46].
Shifting from simulations to the real world, human groups act as teams insofar their members make the effort of reading one other’s mind, predict what their mates will do, and behave accordingly [47] [91]. Experimental tests on indicators of social attentiveness, such as eye movements or taking turns at speaking, confirm this point [100]. Predictions in teams clearly go beyond mere extrapolation, but still, they are not as complex as having an overall picture of how the whole team is working. The coach of a sporting team may have it, but all what is required by team players is that they observe one another, predict, and coordinate.
Out of these considerations we distill the notion that individual ability to make predictions can contribute to collective intelligence. We subsume this concept by means of the following proposition:
Proposition 2.
Individual intelligence, if employed in order to predict and coordinate behavior, can improve collective intelligence.
Finally, let us consider maximum KIDS. While one can think of several fields other than prediction, the ability to think and understand collective behaviour stands out as a clear benchmark for the maximum sophistication of individual agents. This idea can be traced back to Ross Ashby, who called attention onto the difference between the dynamics expressed by the variables of a system and those induced by changing its parameters [5] [84]. At least three influential authors such as Gregory Bateson, Chris Argyris and Donald Schön have been inspired by Ashby when pointing to the importance of understanding system-wide dynamics ([10], p. 292, [3], p. 21, footnote n. 1) [87].
Henceforth, we shall express this idea in terms of Lyapunov functions, saying that while short-sighted decision-makers can be represented by means of variables descending along given basins of attraction, those who are capable to discern collective dynamics are also contemplating the possibility of changing the shape of those basins (see § 3).
Following customary usage in Psychology, let us employ the term meta-cognition for an individual’s capability to understand collective dynamics:
Proposition 3.
Single individuals may be able to envisage the collective intelligence of their organization.
One notable feature of Proposition 3 is that it does not immediately translate into specific consequences for the collective. One critical reason is that, given an individual who is capable of meta-cognition, his possibilities for affecting the organization depend on hierarchical position [66] [49]. Specifically, we singled out three versions of Proposition 3 that differ from one another in this respect.
The first version of Proposition§ 3 is double-loop learning [2]. Known also as “higher-level” [24], “adaptive” [75], “radical” [61], “second-order” [4] or “meta” [1] -learning, it has the action-oriented flavor of management studies where one or a few leaders understand causal relations and conceive innovative solutions. This point of view is entailed in the following version of Proposition 3:
Proposition 3.1.
Single individuals may be able to envisage the collective intelligence of their organization and steer it towards alternative dynamics.
By contrast, if those individuals who envisage the collective intelligence of their organization do not have the authority to steer it, deutero-learning ensues [10] [11]. This concept stems from anthropology and it rather fits the case of individuals who must passively accept a mechanism of which they become aware, whose aims they do not share [86] [88]. Contradiction between official declarations and real experience generates conflicting obligations known as double bind [11] [74]. This point of view is entailed in the following version of Proposition 3:
Proposition 3.2.
Single individuals may be able to envisage the collective intelligence of their organization while not being able to steer it towards alternative dynamics.
Finally, a still different case occurs when the collective intelligence of an organization is envisaged by the vast majority of its stakeholders who openly agree on a course of action [48] [99]. The third version of Proposition 3 ensues:
Proposition 3.3.
The majority of individuals may be able to envisage the collective intelligence of their organization and agree to steer it towards alternative dynamics.
2.1 Technical Qualifications
The emergence of collective dynamics out of individual behaviour has also technical features that are unrelated to the degree of individual intelligence. These technical features can eventually qualify the above propositions.
The first qualification requires individuals to be heterogeneous in order to generate interesting collective dynamics. This qualification applies even to extreme KISS approaches such as artificial neural networks, whose neurons are as simple as summators and yet require random initialization of weights. In more complex models heterogeneity increases the range where self-organized critical states perform the most complex computations [72]. One intuitive confirmation comes from the electrical signals of real brains, whose neurons fire in unison under epileptic seizures but exhibit complex patterns otherwise [12] [38].
Group psychoanalysis offers a similar insight concerning mass behaviour, which can slip into irrational modes where individuals merge their minds generating collectives that do not exhibit any intelligence at all (e.g., certain religious sects, or crowds following populist leaders) [14] [81]. By contrast, crowd wisdom — the ability of large groups to provide solutions that would be difficult for isolated individuals to reach [44] — relies on heterogeneity of individual interpretations [41] [42].
The above considerations can be subsumed in the following qualification:
Qualification 1.
Element heterogeneity is necessary to reach collective intelligence.
The second technical qualification concerns collective agreements and it is obviously most relevant for Proposition § 3.3. Collective agreements generate stable equilibria, but individual intelligence can still make a difference isofar it concerns the speed of convergence.
Macroeconomics has a good point in case. In the 1970s, many Central Banks were trying to stimulate the economy by increasing the money supply in the belief that economic actors would use it to increase GDP, but those actors were eventually able to anticipate that inflation would be the long-term consequence, which expectation generated inflation already in the short run [53]. In this case, greater individual intelligence — labeled as rational expectations — made for faster convergence to a high-inflation, high-unemployment equilibrium. Note that in this case the “agreement” exists only among actors other than the Central Bank (which is treated as an exogenous disturbance, indeed).
Let us capture this insight by means of the following qualification:
Qualification 2.
Given a stable equilibrium generated by a collective agreement, individual intelligence accelerates convergence.
Finally, the third technical qualification concerns the possibility that small individual decisions have a sizeable impact on collectives. Specifically, the butterfly effect states that non-linearities, through chains of bifurcations, are capable of generating chaotic dynamics which may revolve around a “strange attractor” nevertheless [51]. We need a sort of generalized-butterfly qualification that is not limited to chaotic dynamics but includes any sizeable change of the Lyapunov function as well. Let us express this extended butterfly effect by means of the following qualification:
Qualification 3.
Any degree of individual intelligence can trigger non-linear interactions that ultimately generate a substantial impact on collective intelligence.
3 Predators and Preys
In this section we apply the aforementioned insights to the Lotka-Volterra model of prey-predator dynamics [52] [90]. Specifically, we explore whether any sort of collective intelligence appears by endowing predators and preys with variants of individual intelligence inspired by the propositions and qualifications outlined in § 2.
In general, co-existence of predators and preys is the preferred outcome in any application of the Lotka-Volterra model, either because one desires some sort of ecological equilibrium, or because the model performs more sophisticated, more interesting computations if species co-exist, or both. Henceforth, we shall take species co-existence as an indication of collective intelligence, looking for sorts of individual intelligence that make it appear.
There are several reasons for focusing on the Lotka-Volterra model. The first one is that it is relevant to a number of settings other than animal species, including as diverse applications as the business cycle [35] [55] [78] [101], technological substitution [13] [62] [101], ideological and power struggles [89] [17] [59] [40] and market-share dynamics [94] [58] [92].
Thus, it arguably reflects important aspects of human relations and interactions. Exploring the possibility that predators and preys are capable of sophisticated forms intelligence reflects these sorts of applications rather than the purely biological ones, whereas simpler algorithms are likely to be expressed by less sophisticated brains.
The second reason is that although the basic Lotka-Volterra is a mathematically well-known object, recent research has highlighted that its dynamics change substantially if realistic features are added to it, such as random noise [57] [104] [50] [85], time delays [103] [7] [102] and discrete dynamics [70] [68]. Thus, it is a good candidate for ABMs that aim at reconstructing collective dynamics out of individual decisions.
However, existing ABMs achieve consistency with the aggregate Lotka-Volterra by adding features that are specific to animal populations, such as availability of food and shelter for preys, or stages of phenotypic development for predators [97] [98] [96] [15] [83]. By contrast, we are interested in collective dynamics that can be influenced by sophisticated individual decisions. Thus, a final reason is that this endeavour has not been attempted hitherto.
Let us begin with the basic Lotka-Volterra model. Let $x\in\Re^{+}_{0}$ and $y\in\Re^{+}_{0}$ denote the environmental density  [^1] By “environmental density” we mean the ratio of the number of individuals to some measure of the size of the natural environment where predators and preys live, such as its area or volume. of preys and predators, respectively. In its simplest version, the Lotka-Volterra model captures their dynamics by means of the following pair of differential equations:
$\left\{\begin{array}[]{rcl}\dot{x}&=&ax-bxy\\ \dot{y}&=&-cy+dxy\end{array}\right.$ (1)
where $a,b,c,d\in\Re^{+}$ are suitable constants.
The Lotka-Volterra model has three outcomes:
$E_{1}=(0,0)$
No preys, no predators. If $x(0)=0$ and $y(0)=0$, then $x(t)=y(t)0\forall t>0$. However, this equilibrium is also approached if $x(0)=0$ and $y(0)>0$, in which case $x(t)=0\forall t>0$ and $y(t)=y(0)\,e^{-ct}$, which implies that $\lim_{t\to+\infty}y(t)=0$.
$E_{2}=(+\infty,0)$
Only preys, no predators. If $x(0)>0$ and $y(0)=0$, $x(t)=x(0)\,e^{at}$ and $y(t)=0\forall t>0$. The population of preys grows indefinitely because in the limit $\lim_{t\to+\infty}x(t)=+\infty$.
$E_{3}=(c/d,a/b)$
Preys and predators coexist. In the basic Lotka-Voltera, the two species co-exist in a limit cycle where both populations oscillate.
where $E_{*}=(x^{*},y^{*})$ are the equilibrium coordinates.
Let us look for individual behavior that makes $E_{3}$ sustain itself. One possible Lyapunov function for the Lotka-Volterra model is [80]:
$V(x,y)=d(x-c/d\,\ln x)+b(y-a/b\,\ln y)+const$ (2)
with $a$, $b$, $c$ and $b$ as in eq. 1.
It is possible to show that $E_{3}$ appears if coefficients $a$ and $d$ are sufficiently small and coefficients $b$ and $c$ are sufficiently large (see § A). In other words, preys and predators are more likely to co-exist if preys do not grow too quickly and, furthermore, their growth is effectively checked by predators. Predators in their turn should grow with preys, but they must quickly slow down their growth as soon as their population becomes too large.
Let us turn to the discrete-time, agent-based version of the Lotka-Volterra model in order to analyze what behavioral algorithms take the system into either $E_{1}$, $E_{2}$ or $E_{3}$, respectively. Notably, moving onto computational models based on heterogeneous interacting agents implies that Technical Qualification 1 applies.
Henceforth we shall start from Wilensky and Reisman’s Wolf Sheep Predation model [95] [96] on the NetLogo platform where this model is available in the standard library. In order to experiment with behavioral hypotheses we added additional features to the basic model. Our extended version is available on CoMSES.
In Wilensky and Reisman’s Lotka-Volterra, $E_{3}$ almost never sustains itself. Specifically, we observed $E_{1}$ (both preys and predators disappear) $41\%$ of the times, $E_{2}$ (only preys survive) $58.7\%$ of the times whereas only $0.3\%$ of the times the two species managed to co-exist (see Appendix B for details). Wilensky and Reisman obtained stable cycles by adding a renewable source of energy for preys [96].
We took on a different route. We started from the basic, unstable model that almost invariably ends into either $E_{1}$ or $E_{2}$, and experimented with predators’ and preys’ intelligence in order to discover whether either predators’ or preys’ intelligence, or both, can modify its Lyapunov function to make $E_{3}$ the center of a new basin of attraction.
Figure 1 illustrates this concept by means of fictional Lyapunov functions. On the left (a), a saddle from which the system sooner or later ends either in $E_{1}$ or in $E_{2}$. This is the sort of Lyapunov function that likely describes the Wilensky-Reisman model as it is. We would like that preys and predators behave in ways such that the Lyapunov function becomes either as in (b), in which case both species coexist with oscillating populations, or (c), in which case constant populations of predators and preys obtain. Our research question is whether any sort of individual intelligence exists, that helps them reach one of these configurations.
Figure 1: Three Lyapunov functions. Left (a), a saddle with equilibria at its extremes. Center (b), a stable limit cycle along which the two populations oscillate. Right (c), a stable equilibrium point to which the two populations converge. By courtesy of ©Rong Ge.
Henceforth we explore what happens to Wilensky and Reisman’s model by endowing preys and predators with alternative sorts of intelligence inspired by the Propositions and Qualifications expounded in § 2 and § 2.1, respectively. Details about the model are in Appendix B. All reported results have been averaged over 1,000 runs.
3.1 It’s Difficult to Leave a KISS
As illustrated above, the basic Wilensky and Reisman’s model does not support $E_{3}$. Suppose that predators and preys devise a reproductive strategy which, in their intentions, should be able to keep the system at $E_{3}$. With a slight departure from pure KISS, they may imagine that a flexible and adaptive strategy based on some negative feedback may generate co-existence. For instance, in order not to be either too many when the other species shrinks, or too few when the other species thrives, either predators, or preys, or both of them may decide to bind their reproduction to the other species’s success:
  1. 1.
    Predators reproduce proportionally to the fraction of preys.
  2. 2.
    Preys reproduce proportionally to the fraction of predators.
  3. 3.
    Predators reproduce proportionally to the fraction of preys and preys reproduce proportionally to the fraction of predators.
Behaviors 1, 2 and 3 appear sensible but, surprisingly, none of them works very well. With 1 the model ends up in $E_{1}$ $65.3\%$ of the times, in $E_{2}$ $12.8\%$ of the times and only $21.9\%$ of the times in $E_{3}$. With 2 and 3 it is even worse, making the model reach $E_{2}$ $100\%$ of the times (predators go extinct, the population of preys grows indefinitely). One may claim that reproducing proportionately to the fraction of predators is a great strategy for preys because they end up as the sole surviving species independently of what predators do, except for the fact that unlimited growth of one species is not sustainable in the long run.
Thus, this experiment apparently confirms Proposition 1, namely that individual intelligence destroys collective intelligence. However, these results were obtained with individuals who were not even attempting to figure out the global consequences of their behaviour. It was a very slight departure from pure KISS. Perhaps, meta-cognition — the maximum KIDS of Proposition 3 — can help devising individual behavior that generates co-existence.
3.2 The Extreme KIDS of Good Guys
Let us now explore the possibility that, as expressed by Proposition 3, at least some individuals can envision collective dynamics and therefore are able to influence it by selecting some appropriate behavior. Specifically, let us suppose that either predators, or preys, or both of them take on a collaborative attitude that may favor a collective agreement as prescribed by Qualification 3.3.
Let us suppose that the predators realize that, if there are too many of them, preys will disappear so in the end they will have no food and will go extinct. Therefore, they collectively decide to stop reproducing if their population becomes larger than the population of preys because they have an incentive not to endanger their own species’ long-term survival.
Likewise, preys may collectively decide to stop reproducing if their population becomes larger than the population of predators. However, their behavior would be purely altruistic.
Finally, these behaviors can be combined by assuming that both predators and preys stop reproducing if their own population becomes larger than the other one. While the previous decisions were the outcome of agreements between either predators or preys, the combined behavior could be the outcome of a general agreement involving both populations.
  1. 4.
    Predators stop reproducing if their population becomes larger than the population of preys.
  2. 5.
    Preys stop reproducing if their population becomes larger than the population of predators.
  3. 6.
    Predators stop reproducing if their population becomes larger than the population of preys and preys stop reproducing if their population becomes larger than the population of predators.
Note that 4, 5 and 6 are very different from 1, 2 and 3. Since they have been elaborated out of a shared understanding of global dynamics, they non-linearly depend on global thresholds rather than relying on continuous adaptation.
These reproductive strategies are also remarkably similar to those of the mathematical continuous-time model described by eq. 1. Indeed, assuming that predators are willing to stop reproducing if preys are too few (behavior 4) corresponds to a small coefficient $d$ in eq. 1, whereas preys willing to stop reproducing if predators are too few corresponds to a large coefficient $b$. Eventually, small $a$ and large $c$ reinforce the effect of a large $b$ and small $d$.
If predators behave as in 4, $E_{1}$ is reached only $2.8\%$ of the times, $E_{2}$ just $0.5\%$ of the times, whereas $E_{3}$ is reached $96.7\%$ of the times. Thus, this behavior succeeds to make predators and preys co-exist. To a closer scrutiny, this $96.7\%$ arises out of a $43.7\%$ of outcomes where the two populations are roughly constant, a $13.2\%$ where the population of preys oscillates whereas predators do not, a $21.0\%$ where the population of predators oscillates whereas preys do not and a $18.8\%$ where both populations oscillate.
However, if preys behave as in 5 the outcome is radically different. In this case, $E_{1}$ (extinction of both predators and preys) occurs $0.1\%$ of the times, $E_{2}$ (exclusive survival and indefinite growth of preys) happens an overwhelmingly $99.8\%$ of the times whereas with a mere $0.1\%$ $E_{3}$ (co-existence) is just as unlikely as generalized extinction. Once again, preys’ altruistic behavior turns to their own advantage in the short run, but an ever-increasing population of preys is not sustainable in the long run. Once again, collective intelligence is not there.
If both predators and preys take on the collaborative attitude 6 the outcome is somewhat mixed, with $E_{1}$ nonexistent at $0.0\%$ whereas both $E_{2}$ and $E_{3}$ are quite substantial at $22.9\%$ and $77.1\%$, respectively. Co-existence is the most likely outcome but happens less often than in 4. Unlimited growth of preys is substantial but far from inevitable as in case 5.
Here it is worth noticing that, in the Lotka-Volterra model, predators and preys do not enjoy the same status. On the one hand, predators capable of meta-cognition are in position to steer the ecology towards co-existence by exerting double-loop learning (Proposition 3.1). On the other hand, preys capable of meta-cognition understand the system but cannot steer it, as it is typical of deutero-learning (Proposition 3.2). Preys can at most obtain dominance in the short run, but at the cost of unsustainable explosive dynamics for themselves.
One may remark that case 4 is remindful of the sort of relations that humans (the predators) are entertaining with the ecosystem (the preys). Humans are predators who are capable of meta-cognition, and therefore they are in a position that enables them to steer the ecosystem towards sustainable co-existence. By contrast, cases 5 and 6 do not appear to have an immediate counterpart in the real world.
3.3 The Extreme KIDS of Realpolitik
The problem with behavioral rules 4, 5 and 6 is that either preys or predators, or both, are supposed to be capable of sacrificing their immediate interests. In practice, this may be difficult to attain.
Is it possible to obtain stable coexistence of species without such a high degree of public spirit? We still want individuals to use their intelligence to understand the system’s global dynamics, but we also want them to pursue their own interests.
Proposition 2 can be combined with Proposition 3 to endow our agents with the ability to make predictions based on global understanding of the Lotka-Volterra model that they inhabit. Specifically, predators predict their own extinction if they either see the population of preys or their own population growing very quickly, whereas preys predict their own extinction if they see predators growing very quickly.
If they are rational, neither predators nor preys wait for their Lotka-Volterra world to unfold their inevitable destiny. As Qualification 2 suggests, they collectively agree to commit suicide if they see their inevitable end forthcoming. Thus, let us explore the following behavioural rules:
  1. 7.
    Predators commit suicide if either the population of preys is growing much faster than their own population, or if their own population is growing much faster than the population of preys.
  2. 8.
    Preys commit suicide if the population of predators is growing much faster than their own population.
  3. 9.
    Predators commit suicide if either the population of preys is growing much faster than their own population, or if their own population is growing much faster than the population of preys, and preys commit suicide if the population of predators is growing much faster than their own population.
These behavioural rules require an additional parameter to specify what “much faster” means. We introduced this parameter with the default assumption that “much faster” means ten times faster, supporting this assumption with a sensitivity analysis reported in Appendix C.
The results are quite discomforting. With 7, $E_{2}$ is reached $100\%$ of the times. With either 8 or 9, $E_{1}$ (generalized extinction) is reached $100\%$ of the times.
Notably, by assuming that either predators or preys, or both of them are capable to predict the correct and inevitable outcome of the Lotka-Volterra model, co-existence $E_{3}$ can never be reached. Apparently, one cannot escape KISS (Proposition 1).
3.4 Prediction based on Extrapolation
Let us consider the possibility that predators and preys make predictions by making use of linear extrapolation, which amounts to combining Proposition 2 with some intermediate level between KISS (Proposition 1) and KIDS (Proposition 3). In particular, let us assume that either predators, or preys, or both of them reproduce out of extrapolation of the other species’ reproductive trends:
  1. 10.
    Predators reproduce with probability proportional to the variation of the number of preys in the last simulation steps.
  2. 11.
    Preys reproduce with probability proportional to the variation of the number of predators in the last simulation steps.
  3. 12.
    Predators reproduce with probability proportional to the variation of the number of preys in the last simulation steps and preys reproduce proportionally to the variation of the number of predators in the last simulation steps.
Similarly to§ 3.3, also in this case it is necessary to add a parameter in order to specify how many past steps the above variations are computed on. We used this parameter with the default value of five steps, supporting this choice with a sensitivity analysis reported in Appendix C.
Behaviour 10 is the second most effective in generating co-existence after 4 (altruistic predators seeking a collective agreement, § 3.2). In particular, with 10 generalized extinction $E_{1}$ never occurs. Preys-dominated equilibrium $E_{2}$ is reached $8.4\%$ of the times and, most importantly, $E_{3}$ is reached $91.6\%$ of the times. Out of this $91.6\%$, in $7.7\%$ both populations oscillate, $78.1\%$ only predators oscillate, in just $0.4\%$ of cases the opposite happens and, finally, in the remaining $5.4\%$ neither population oscillates. In a nutshell, the most likely outcome is that the population of predators oscillates while the population of preys remains roughly constant.
By contrast, in the opposite case 11 generalized extinction $E_{1}$ is reached $99.5\%$ of the times, whereas in the remaining $0.5\%$ $E_{2}$ is reached. Most notably, $E_{3}$ is never reached.
Once again, in the Lotka-Volterra model there exists a clear asymmetry between predators and preys insofar it concerns what they can achieve with any given level of individual intelligence. Specifically, intelligent predators can attain co-existence by making predictions (Proposition 3.1), whereas equally intelligent preys cannot (Proposition 3.2).
Interestingly, the mixed arrangement 12 is not some sort of average between 10 and 11 but rather a case on its own. With 12, $E_{1}$ is reached $17.9\%$ of the times, $E_{2}$ a mere $2.4\%$ whereas $E_{3}$ is reached $79.7\%$ of the times. Out of this $79,7\%$, $61.7\%$ of the times both populations oscillate, $1.4\%$ only the population of predators oscillates, $14.2\%$ the opposite happens and, finally, in the remaining $2.4\%$ neither population oscillates. Thus, co-existence is reached quite often, and most of the times it is reached with both populations oscillating.
This sort of co-existence is quite remarkable because, differently from all previous cases where $E_{3}$ occurred, both populations explode. A large herd of preys moves in our artificial space, hunted by an even larger herd of predators. Preys reproduce while running away from predators, predators reproduce while chasing preys generating the two moving columns illustrate in Figure 2.
Figure 2: The most common outcome of case 12. Both populations grow while predators (black wolves) are chasing preys (white sheep).
Figure 3 illustrates the average of the 797 time series that reach co-existence ($E_{3}$), plotted against one of them in order to show individual variability. Both populations oscillate, with the population of predators growing faster than preys. Thus, while in all previous cases co-existence was associated with sustainability, in case 12 co-existence of predators and preys does not imply a sustainable future for either species.
Figure 3: The oscillations of the population of predators, averaged over the 797 time series where $E_{3}$ obtains (thick red line) plotted against one single time series selected for being closest to the mean (thin black line). The oscillations of the population of preys, averaged over the 797 time series where $E_{3}$ obtains (thick blue line) plotted against one single time series selected for being closest to the mean (gray thin line). On average, the population of predators grows from 50 to 49,530.68 individuals whereas the population of preys grows from 100 to 7,453.59 individuals.
One may argue that humans are the only species able to make predictions based on variations of the populations of other species, and that since humans are predators, only the case 10 occurs in the real world. Specifically, humans can achieve co-existence with other species by observing their variations and behaving accordingly. This is interesting in itself, but we already mentioned that the Lotka-Volterra model has many interpretations besides the purely ecological one.
In particular, let us consider the application of the Lotka-Volterra model to the rather irregular oscillations of economic activities with a period of about 8-10 years, known as business cycle or trade cycle [35] [55] [78]. This tradition goes back to Marx, who remarked that capitalists cannot make profits if after many years of economic growth unemployment has become so low and salaries so high that capitalists do not make a profit by running their companies. This circumstance plunges the economy into recession, until unemployment becomes sufficiently high to induce workers to accept lower salaries again [60]. The Lotka-Volterra model can capture this dynamics by identifying capitalists with predators and workers with preys [35] [55] [78].
However, the aggregate Lotka-Volterra generates business cycles without economic growth. By contrast, our agent-based Lotka-Volterra combines growth with the business cycle. This aspect of our model deserves some comments.
Capitalism is — among else — a way of thinking, based on postponing consumption in order to save and invest. It requires making extrapolations about the future, a circumstance which, according to our computational experiments, turns oscillations around a fixed point into oscillations around a growing average. Notably, our model suggests that economic growth can only set in once the vast majority of actors have acquired an extrapolation-based way of thinking, a circumstance which in pre-modern societies may have never realized. Albeit kings did order the construction of infrastructures that would benefit society at large in the long run, insofar most people reasoned in terms of enjoying life day-by-day (carpe diem), indefinite economic growth could not start.
One interesting question is whether the explosive oscillations originated by behavior 12 are an instance of the extended butterfly effect (Technical Qualification 3). Do economic actors want indefinite growth, or do they want equilibrium with the rest of the ecosystem? How much aware are they of the collective intelligence of the system as a whole? Qualification 3 suggests that indefinite growth may be an emergent property, well beyond individual decisions.
4 Conclusions
Since we investigated such ubiquitous dynamics as those expressed by preys and predators, one is tempted to ask whether our findings have some validity beyond this specific model. In particular, one may sensibly ask whether there exists a wider class of models where individual prediction by linear extrapolation generates interesting collective dynamics. And, one may also extend this question from the first to the second derivative and so forth, asking whether higher-order derivatives at the individual level are key to generate novel collective dynamics.
Tentatively, we are inclined to answer the first question with a “Yes” upon condition that the vast majority of individuals make predictions, whereas we bet that the answer to the second question is a clear “No.” Our rationale is that if most individuals compute first-order derivatives, then they can impact on one another generating higher-order derivatives. This is likely to happen in any model where interactions generate multiplicative, snowball-like effects that end up with constructing higher-order derivatives at the collective level. By contrast, higher-order derivatives at the individual level would add very little to the possibilities for interesting collective dynamics and, if anything, they could generate overfitting and individual complex dynamics cancelling out. These are pure speculations but, in a nutshell, the idea is that distributed systems are capable of doing sophisticated computations if individuals provide them with the basic bricks, namely, first-order derivatives.
The discovery of a coexistence-with-explosion dynamics that emerges in the prey-predator model if all agents make linear predictions highlights the presence of conflicting values in science, as well as in society at large. While everybody agrees that co-existence of species is a good thing, explosive dynamics is generally valued very differently by economists (GDP growth) and environmentalists (non-sustainable anthropic activities).
Evidently, the difference is in implicit predictions. Environmentalists implicitly predict that explosive dynamics will eventually bring about generalized extinction, whereas the economists do not. Both positions are tenable, precisely because they are implicit. Since this sort of long-run dynamics are not in the model, whether co-existence plus explosive dynamics represents an instance of collective intelligence is, within this model, an undecidable question.
Conflicts of interest and other legal disclaimers
The authors declare not to have any conflict of interest whatsoever with the organizations, institutions and persons directly or indirectly mentioned in this paper. The authors did not receive any funding to carry out this research.
The authors received permission to reproduce the images that appear in Figure 1 from their copyright holders. All other figures and tables have been generated by the authors. No empirical data have been used.
Authors are listed in alphabetical order. GF contributed the theoretical architecture and literature review, AP contributed the computational model and exploration of results. The behavioural choices of artificial agents have been thoroughly discussed by both authors.
Appendix A Lyapunov’s Stability Theorem and its Application to the Lotka-Volterra Model
Henceforth, Lyapunov’s Stability Theorem will be illustrated for two-dimensional systems. Subsequently, it will be applied to the Lotka-Volterra prey-predator model by developing a Lyapunov function specifically designed to highlight the influence of the coefficients of eqs. (1) on the stability of the limit cycle.
Consider a nonlinear two-dimensional dynamical system:
$\left\{\begin{array}[]{rcl}\dot{x}&=&f(x,y)\\ \dot{y}&=&g(x,y)\end{array}\right.$
where $x,y\in\Re$ and $f,g\in\Re^{2}\mapsto\Re$.
Without any loss of generality, let us assume that this system has an equilibrium point at $E=(x^{*},y^{*})$. We want to know whether this equilibrium is stable.
The Lyapunov Theorem states that if a function $V:\Re^{2}\mapsto\Re$ exists, such that:
  1. 1.
    $V(x,y)=0$ if and only if $(x,y)=(x^{*},y^{*})$;
  2. 2.
    $V(x,y)>0$ if and only if $(x,y)\neq(x^{*},y^{*})$;
  3. 3.
    $\dot{V}(x,y)=\frac{d}{dt}V(x,y)=\frac{\partial V}{\partial x}f(x,y)+\frac{\partial V}{\partial y}g(x,y)\leq 0$ for $(x,y)\neq(x^{*},y^{*})$;
then $(x^{*},y^{*})$ is a stable equilibrium point. Function $V(x,y)$ is called a Lyapunov function.
The Lotka-Volterra model (1) has an equilibrium point at $(x^{*},y^{*})=(c/d,\,a/b)$. Let us consider the following function:
$V(x,y)=d(x-x^{*}\ln x)+b(y-y^{*}\ln y)+[d(x^{*}-x^{*}\ln x^{*})+b(y^{*}-y^{*}\ln y^{*})]$
which we may also write as $V(x(t),y(t))=d(x-c/d\ln x)+b(y-a/b\ln y)+const$.
It is obviously $V(x^{*},y^{*})=0$. Thus, condition (1) is satisfied.
Let us check whether $V(x,y)>0$. Let us consider the first term, namely, $d(x-x^{*}\ln x)$. It is $d(x-x^{*}\ln x)>0$ if $dx-c\ln x>0$rightarrow $dx>c\ln x$rightarrow $x>c/d\ln x$rightarrow $e^{x}>e^{c/d\ln x}$rightarrow $e^{x}>x^{c/d}$ which is always true because the exponential function yields greater values than the power function. The third term, $d(x^{*}-x^{*}\ln x^{*})$, makes sure that the combination of the first and the third term starts to yield positive values just as soon as $V$ leaves equilibrium $(x^{*},y^{*})$. Likewise, $e^{y}>y^{a/b}$ is always true because the exponential function yields greater values than the power function and the fourth term $b(y^{*}-y^{*}\ln y^{*})$ ensures that this only happens outside $(x^{*},y^{*})$. Thus, condition (2) is satisfied.
Let us turn to the third condition. It is $\dot{V}$ $=$ $(d-d\frac{x^{*}}{x})(ax-bxy)+(b-b\frac{y^{*}}{y})(-cy+dxy)$ $=$ $ad(x-x^{*})-bc(y-y^{*})+bd(x^{*}y-y^{*}x)$. Obviously, $\dot{V}(x^{*},y^{*})=0$. Let us investigate which parameters make for $\dot{V}(x,y)<0$ at $(x,y)\neq(x^{*},y^{*})$.
Let us explore the sign of $\partial\dot{V}/\partial x=ad-bdy^{*}$ and $\partial\dot{V}/\partial y=-bc+bdx^{*}$. $\dot{V}$ is negative if:
$\left\{\begin{array}[]{rcl}ad-bdy^{*}&<&0\\ -bc+bdx^{*}&<&0\end{array}\right.$
We obtain that $\dot{V}$ is negative if:
$\left\{\begin{array}[]{rcl}a&<&by^{*}\\ c&>&dx^{*}\end{array}\right.$
This means that:
  • The smaller $a$, the more likely that $\dot{V}<0$;
  • The greater $b$, the more likely that $\dot{V}<0$.
  • The greater $c$, the more likely that $\dot{V}<0$.
  • The smaller $d$, the more likely that $\dot{V}<0$.
Thus, we can conclude that with small $a$, large $b$, large $c$ and small $d$ the equilibrium $(c/d,a/b)$ is more likely to be stable.
Appendix B Our Agent-Based Lotka-Volterra
Since we built our model out of Wilensky and Reisman’s model [95] [96], our code keeps naming predators as “Wolves” and preys as “Sheep,” respectively. Specifically, we derived our model from the sheep-wolves version which lacks the agent “Grass.” Correspondingly, we eliminated the parameters grass-regrowth-time, sheep-gain-from-food. By contrast, parameters initial-sheep, initial-wolves, sheep-reproduce, wolf-reproduce and wolves-gain-from-food carry on to our model.
We kept the five parameters that we mutuated from Wilensky and Reisman’s model at their original default values, whereas we used switches to select among behavioral configurations in § 3. We introduced two new parameters in order to explore the behavior configurations of § 3.3 and § 3.3, which we named how-much-faster and Timespan, respectively. We kept these two additional parameters at their base values throughout our simulations, carrying out a sensitivity analysis whose results are reported in Appendix C. The ensuing Table 1 summarizes the parameter values of our model for each behavior configuration, as well as the switches that we used to select among them.
Behavior Configuration
Parameter Name 0 1 2 3 4 5 6 7 8 9 10 11 12
initial-sheep 100 100 100 100 100 100 100 100 100 100 100 100 100
initial-wolves 50 50 50 50 50 50 50 50 50 50 50 50 50
sheep-reproduce 4 4 4 4 4 4 4 4 4 4 4 4 4
wolf-reproduce 5 5 5 5 5 5 5 5 5 5 5 5 5
wolves-gain-from-food 20 20 20 20 20 20 20 20 20 20 20 20 20
Wolves-repr-by-frac-sheep Off On Off On Off Off Off Off Off Off Off Off Off
Sheep-repr-by-frac-wolves Off Off On On Off Off Off Off Off Off Off Off Off
Stop-rep-if-wolves$>$sheep Off Off Off Off On Off On Off Off Off Off Off Off
Stop-rep-if-sheep$>$wolves Off Off Off Off Off On On Off Off Off Off Off Off
Wolves-perfect-foresight Off Off Off Off Off Off Off On Off On Off Off Off
Sheep-perfect-foresight Off Off Off Off Off Off Off Off On On Off Off Off
how-much-faster 10 10 10 10 10 10 10 10 10 10 10 10 10
Wolves-extrapolate Off Off Off Off Off Off Off Off Off Off On Off On
Sheep-extrapolate Off Off Off Off Off Off Off Off Off Off Off On On
Timespan 5 5 5 5 5 5 5 5 5 5 5 5 5
Simulation Length 500 500 500 500 500 500 500 500 500 500 500 500 300
max-wolves 500k 500k 500k 500k 500k 500k 500k 500k 500k 500k 500k 500k   1M
max-sheep 500k 500k 500k 500k 500k 500k 500k 500k 500k 500K 500k 500k   1M
Transitory 50 50 50 50 50 50 50 50 50 50 50 50 50
Benchmarking 100 100 100 100 100 100 100 100 100 100 100 100 100
Table 1: The switches and parameters of our model, for each behavior configuration. Switches that are “On” for a specific configuration are highlighted in yellow. The basic Wilensky model is denoted as the zero-configuration, the remaining ones are numbered as in § 3. In the top five lines, the five parameters inherited from Wilensky’s model. Subsequently, the two parameters relative to configurations 1, 2, 3 and the two parameters relative to configurations 4, 5, 6, respectively. Then the three parameters that are relative to configurations 7, 8, 9 and 10, 11, 12, respectively. In the bottom lines, the five parameters that have been introduced in order to run and observe the model.
By means of the parameters Simulation Length, max-wolves and max-sheep our model runs until either the maximum number of steps is reached, or both wolves and sheep have died out, or either population has reached the maximum allowed. In most behavioral configurations we ran the model for a maximum of 500 steps and allowed a maximum of 500,000 wolves and 500,000 sheep, respectively. These values were chosen because 500 steps were more than enough to observe results, and because the 500,000 threshold was reached only when either wolves or sheep were the only surviving species. However, case 12 was quite different because both populations were growing very fast and, specifically, the population of wolves was growing too large for available computational resources. In order to minimize the number of times when the population threshold would be reached we increased it from 500,000 to 1,000,000 individuals and, since with this behavioral configuration the pattern became very clear very early, we shortened the simulation length from 500 to 300 steps. Even with these values, the population of wolves hit the threshold in a $7.9\%$ of our 1,000 runs.
In order to observe outputs we also introduced a parameter Transitory and a parameter Benchmarking that we set at 50 and 100 steps, respectively. Between Transitory and Benchmarking the model determines which dynamics count as oscillations. At each step within this interval, the minimum and maximum number of sheep and wolves that have been ever attained are updated. Let us call them $MINs$, $MAXS$, $MINW$ AND $MAXW$, respectively. Sheep oscillations are counted if $(MAXS-MINS)/2\geq 0.2(MAXS+MINS)/2$. If this condition is satisfied, one oscillation occurs if the number of sheep is greater or equal to $(MAXS+MINS)/2+0.5(MAXS-MINS)/2)$ and, subsequently, the number of wolves is smaller or equal to $(MAXW+MINW)/2)-0.5(MAXW-MINW)/2$. Wolves oscillations are counted in a similar way. Also the possibility of perfect foresight is only considered after Benchmarking steps have elapsed.
The parameters that have been introduced in order to run the model and obtain outputs are listed at the bottom of Table 1. Since we are dealing with a simulation model that must be necessarily stopped after a finite number of steps, we eventually obtain outcomes that exhibit richer details than the mathematically defined $E_{1}$, $E_{2}$ and $E_{3}$. We mapped simulation outcomes onto $E_{1}$, $E_{2}$ and $E_{3}$ as follows:
  1. 0.
    When the simulation is stopped, neither sheep nor wolves exist. This case corresponds to $E_{1}$.
  2. 1.
    When the simulation is stopped, only wolves exist. Since this is untenable in the long run, we assumed that wolves would go extinct as well. Thus, also this case corresponds to $E_{1}$.
  3. 2.
    When the simulation is stopped, only sheep exist. This case corresponds to $E_{2}$.
  4. 3.
    Both sheep and wolves exist when the simulation is stopped, and no oscillation has been observed. We classified this case as $E_{3}$.
  5. 4.
    Both sheep and wolves exist when the simulation is stopped. The number of sheep has been oscillating, but no oscillation of the number of wolves has been observed. We classified this case as $E_{3}$.
  6. 5.
    Both sheep and wolves exist when the simulation is stopped. The number of wolves has been oscillating, but no oscillation of the number of sheep has been observed. We classified this case as $E_{3}$.
  7. 6.
    Both sheep and wolves exist when the simulation is stopped. Both the number of sheep and the number of wolves have been oscillating. We classified this case as $E_{3}$.
  8. 7.
    The simulation is stopped because the number of sheep reached a threshold imposed by available computational power. When the simulation was stopped the population of sheep was growing faster than the population of wolves. By extrapolating a future where sheep continue to grow whereas wolves become negligible and finally die out we classified this case as $E_{2}$.
  9. 8.
    The simulation is stopped because the number of sheep reached a threshold imposed by available computational power. Contrary to the previous case, when the simulation was stopped the population of wolves was growing faster than the population of sheep. By extrapolating a future where wolves dominate the scene, eat all the sheep and finally go extinct we classified this case as $E_{1}$.
  10. 9.
    The simulation is stopped because the number of wolves reached a threshold imposed by available computational power. When the simulation was stopped the population of wolves was growing faster than the population of sheep. By extrapolating a future where wolves dominate the scene, eat all the sheep and finally go extinct we classified this case as $E_{1}$.
  11. 10.
    The simulation is stopped because the number of wolves reached a threshold imposed by available computational power. Contrary to the previous case, when the simulation was stopped the population of sheep was growing faster than the population of wolves. By extrapolating a future where sheep continue to grow whereas wolves become negligible and finally die out we classified this case as $E_{2}$.
Our extended version of the Lotka-Volterra model is available on CoMSES. Further documentation is provided along with the code.
Appendix C Sensitivity Analysis
We carried out a sensitivity analysis on the two parameters that we added to Wilensky’s model [95] [96], namely how-much-faster and Timespan. Since how-much-faster regulates the difference of growth rates between wolves and sheep that activates their perfect foresight, we analyzed its impact when behavioral rules (7), (8) and (9) are adopted. Likewise, since Timespan regulates the time interval over which extrapolations are made, we analyzed its impact when behavioral rules 10, 11 and 12 are adopted.
In both cases we explored the consequences on $E_{1}$, $E_{2}$ and $E_{3}$ of $40\%$ increments and decrements of these parameters with respect to their base values. For how-much-faster this meant exploring the consequences of decreasing it to 6 and increasing it to 14 from its base value 10. For Timespan this meant exploring the consequences of decreasing it to 3 and increasing it to 7 from its base value 5. In both cases we measured the occurrence of $E_{1}$, $E_{2}$ and $E_{3}$ over 1,000 runs.
Table 2 reports the outcomes of the sensitivity analysis for parameter how-much-faster. The outcomes with this parameter at its base value are reported in the three central columns, whereas the outcomes when this parameter is decreased by $40\%$ or increased by $40\%$ are reported in the three columns on their left and their right, respectively. It appears that $40\%$ variations of this parameter have no impact whatsoever.
how
muc 6 10 14
fast
be
hav 7 8 9 7 8 9 7 8 9
ior
$E_{1}$ 0 1,000 1,000 0 1,000 1,000 0 1,000 1,000
$E_{2}$ 1,000 0 0 1,000 0 0 1,000 0 0
$E_{3}$ 0 0 0 0 0 0 0 0 0
Tot. 1,000 1,000 1,000 1,000 1,000 1,000 1,000 1,000 1,000
Table 2: The number of aggregate outcomes $E_{1}$, $E_{2}$ and $E_{3}$ with behavioral rules 7, 8 and 9 when how-much-faster is decreased by $40\%$ from 10 to 6 and increased by $40\%$ from 10 to 14, respectively. All other parameters at their base values.
Table 3 reports the outcomes of the sensitivity analysis for parameter Timespan. The three central columns report outcomes with this parameter at its base value, whereas the three columns on their left and their right report the outcomes when this parameter is decreased by $40\%$ and increased by $40\%$, respectively.
It appears that behavioral rule 10 — predators extrapolate, preys do not — is the most sensitive to this parameter. In particular, with $\mathbf{Timespan}=7$ the prevalence of $E_{3}$ disappears, with $E_{2}$ and $E_{3}$ becoming almost equally likely.
By contrast, behavioral rule 11 — preys extrapolate, predators do not — is only marginally affected by Timespan. In spite of $E_{2}$ and $E_{3}$ appearing or disappearing across the three values of Timespan, $E_{1}$ remains the most likely outcome by far.
Finally, behavioral rule 12 — both predators and preys extrapolate — is somehow in between. Similarly to behavior 10, also in this case $\mathbf{Timespan}=7$ has an impact, not strong enough to destroy the overall pattern but sufficient to make $E_{1}$ substantially more likely, $E_{3}$ substantially less likely with respect to the base case.
Ti-
me 3 5 7
sp.
be
hav 10 11 12 10 11 12 10 11 12
ior
$E_{1}$ 0 990 177 0 998 179 0 984 332
$E_{2}$ 1 0 4 91 2 24 457 16 17
$E_{3}$ 999 10 819 909 0 797 543 0 651
Tot. 1,000 1,000 1,000 1,000 1,000 1,000 1,000 1,000 1,000
Table 3: The number of aggregate outcomes $E_{1}$, $E_{2}$ and $E_{3}$ with behavioral rules (10), 11 and (12) when Timespan is decreased by $40\%$ from 5 to 3 and increased by $40\%$ from 5 to 7, respectively. All other parameters at their base values.